Assume Young's modulus for bone is 1.5 x 1010N/m². The bone breaks if stress greater than 1.5 x 108 N/m² is imposed on it. 1) What is the maximum force that can be exerted on the femur bone in the leg if it has a minimum effective diameter of 2.5 cm? 2) If this much force is applied compressively, by how much does the 25 cm long bone shorten?


1
Expert's answer
2021-09-30T14:29:41-0400

Pressure(P)="\frac{F}{A}"

"F=P\times A"

Area

"A={\pi }\times{\frac{2.5^2\times10^{-4}}{4}}"


"F=1.5\times10^{8}\times\pi \times\frac{2.5^2\times10^{-4}}{4}=7.35\times10^4N"

We know that

"Y=\frac{P}{\frac{∆L}{L}}"

"∆L=\frac{P\times L}{Y}"

Put value

"∆L=\frac{1.5\times10^{8}\times 0.25}{1.5\times10^{10}}=0.0025m=2.5mm"